# DPNet: Doppler LiDAR Motion Planning for Highly-Dynamic Environments

### v1

@@ 10 Hz 5 Hz 2 Hz
KF [jian2023dynamic] -22.66 $\pm$ 6.41 -16.62 $\pm$ 5.24 -7.18 $\pm$ 3.84
KNet [revach2022kalmannet] -23.38 $\pm$ 7.82 -17.79 $\pm$ 7.77 -12.56 $\pm$ 7.16
D-KF [peng2021detection] -27.15 $\pm$ 6.39 -21.66 $\pm$ 4.77 -14.09 $\pm$ 3.60
@@
KF [jian2023dynamic] -27.15 $\pm$ 6.35 -20.48 $\pm$ 4.79 -9.15 $\pm$ 6.64
KNet [revach2022kalmannet] -29.26 $\pm$ 5.47 -23.66 $\pm$ 5.98 -17.38 $\pm$ 7.36
D-KF [peng2021detection] -32.74 $\pm$ 6.18 -27.60 $\pm$ 5.11 -19.20 $\pm$ 5.03
@@
KF [jian2023dynamic] -18.50 $\pm$ 6.53 -11.49 $\pm$ 5.92 1.88 $\pm$ 6.47
KNet [revach2022kalmannet] -17.50 $\pm$ 7.49 -11.83 $\pm$ 7.64 -6.32 $\pm$ 6.73
D-KF [peng2021detection] -23.68 $\pm$ 6.44 -17.36 $\pm$ 4.65 -6.61 $\pm$ 4.92
@@
KF [jian2023dynamic] -22.55 $\pm$ 6.46 -14.14 $\pm$ 5.27 8.05 $\pm$ 0.30
KNet [revach2022kalmannet] -23.19 $\pm$ 5.54 -17.34 $\pm$ 6.12 -9.81 $\pm$ 8.92
D-KF [peng2021detection] -29.67 $\pm$ 6.90 -22.97 $\pm$ 5.59 -10.59 $\pm$ 7.26
@@
TABLE II: NMSE (dB) of motion prediction: mean $\pm$ std across tracking frequencies and scenarios on the AevaScenes [aevascenes] dataset. Purple denotes the best mean.
Qualitative Analysis. Fig. 7 demonstrates a visualized result of tracking one rapidly approaching vehicle (at about 25 m/s) on the Highway scenario. Among the evaluated methods, D-KNet demonstrates superior performance by effectively leveraging Doppler cues to enable accurate, real-time tracking of rapid objects, achieving the lowest tracking error of 21.18–21.18 dB. In contrast, KF and KNet struggle to handle the approaching vehicle with abrupt accelerations. This comparison highlights the importance of integrating model-based learning and Doppler LiDAR measurements for more robust tracking under highly-dynamic conditions.
Quantitative Analysis. Table II reports the results across tracking frequencies (2 - 10 Hz), prediction horizons ($\mathcal{H}_{t}=5,10$), and driving scenarios (Highway, City). Across all settings, D-KNet consistently achieves the lowest NMSE and significantly outperforms all baselines. Notably, at $\mathcal{H}_{t}=5$, D-KNet attains $45.00\pm 7.35$ dB in the City scenario at 10 Hz and surpasses the second best D-KF [peng2021detection] by $12.26$ dB. Even at $\mathcal{H}_{t}=10$, D-KNet maintains robustness to long-horizon uncertainty, whereas other methods suffer considerable degradation. Tighter standard deviations of D-KNet further confirm its stability across diverse obstacle dynamics. These results demonstrate the superior reliability of D-KNet, especially under real-world constraints such as varying sensor noise, limited processing frequency, and scenario complexity.
@@ V-C More Quantitative Results
TABLE III: D-KNet’s hardware efficiency across tracked obstacle numbers on NVIDIA Jetson Orin NX 16 GB.
@@
KF [jian2023dynamic] -15.53 $\pm$ 8.80 -15.15 $\pm$ 9.63 -14.72 $\pm$ 8.45 -14.24 $\pm$ 8.26 -13.73 $\pm$ 8.09 -13.20 $\pm$ 7.90 -12.66 $\pm$ 7.72 -12.08 $\pm$ 7.53 -11.52 $\pm$ 7.33 -10.96 $\pm$ 7.15
KNet [revach2022kalmannet] -30.82 $\pm$ 3.80 -24.80 $\pm$ 4.05 -21.44 $\pm$ 4.28 -18.93 $\pm$ 4.50 -16.93 $\pm$ 4.70 -15.24 $\pm$ 4.87 -13.78 $\pm$ 5.03 -12.42 $\pm$ 5.19 -11.23 $\pm$ 5.34 -10.16 $\pm$ 5.50
@@
KF [jian2023dynamic] -16.49 $\pm$ 7.12 -16.34 $\pm$ 6.83 -16.32 $\pm$ 6.69 -16.15 $\pm$ 6.54 -15.92 $\pm$ 6.42 -15.65 $\pm$ 6.30 -15.34 $\pm$ 6.17 -14.98 $\pm$ 6.04 -14.59 $\pm$ 5.91 -14.16 $\pm$ 5.77
KNet [revach2022kalmannet] -38.22 $\pm$ 3.05 -32.42 $\pm$ 3.10 -28.90 $\pm$ 3.16 -26.26 $\pm$ 3.42 -24.13 $\pm$ 3.57 -22.3 $\pm$ 3.72 -20.68 $\pm$ 3.88 -19.22 $\pm$ 3.96 -17.87 $\pm$ 3.95 -16.59 $\pm$ 4.06
D-KF [peng2021detection] -16.29 $\pm$ 8.89 -16.26 $\pm$ 8.19 -16.23 $\pm$ 7.78 -16.19 $\pm$ 7.47 -16.17 $\pm$ 7.26 -16.13 $\pm$ 7.07 -16.13 $\pm$ 6.88 -16.13 $\pm$ 6.68 -16.12 $\pm$ 6.51 -16.11 $\pm$ 6.34
@@
TABLE IV: NMSE (dB) of per-horizon-step motion prediction for $\mathcal{H}_{t}=10$: mean $\pm$ std across scenarios on the AevaScenes [aevascenes] dataset. Purple denotes the best mean.
We also deploy D-KNet on an NVIDIA Jetson Orin NX 16 GB to evaluate its real-time tracking efficiency on a resource-constrained edge platform. It is found that D-KNet consumes only about $107~\text{MB}$ of GPU memory. As seen from Table III, it achieves inference rates of over $100$ Hz when tracking a single fast-moving obstacle. Notably, even with $10$ simultaneously tracked obstacles, D-KNet is able to sustain about $15$ Hz and a low CPU usage of $42.0\,\%$. These results confirm D-KNet’s suitability for lightweight deployment in real-time autonomous systems.
Table IV shows per-step NMSE (dB) over a 10-step horizon. D-KNet consistently outperforms all baselines at every step, with margin widening especially in short steps (e.g., 48.81-48.81 dB of DKNet and 16.49-16.49 dB of KF for step $1$ in City). The KF and D-KF methods exhibit nearly flat error growth, indicating trivial estimates under transition model mismatch. However, D-KNet and KNet show physically reasonable error accumulation, demonstrating the effectiveness of model-based learning against dynamics uncertainty. Notably, D-KNet maintains tighter standard deviations and significantly lower mean error, verifying that the Doppler LiDAR integration enhances both accuracy and robustness for long-horizon predictions.

### v2

@@ 10 Hz 5 Hz 2 Hz
KF [10] -22.66 $\pm$ 6.41 -16.62 $\pm$ 5.24 -7.18 $\pm$ 3.84
KNet [17] -23.38 $\pm$ 7.82 -17.79 $\pm$ 7.77 -12.56 $\pm$ 7.16
D-KF [16] -27.15 $\pm$ 6.39 -21.66 $\pm$ 4.77 -14.09 $\pm$ 3.60
@@
KF [10] -27.15 $\pm$ 6.35 -20.48 $\pm$ 4.79 -9.15 $\pm$ 6.64
KNet [17] -29.26 $\pm$ 5.47 -23.66 $\pm$ 5.98 -17.38 $\pm$ 7.36
D-KF [16] -32.74 $\pm$ 6.18 -27.60 $\pm$ 5.11 -19.20 $\pm$ 5.03
@@
KF [10] -18.50 $\pm$ 6.53 -11.49 $\pm$ 5.92 1.88 $\pm$ 6.47
KNet [17] -17.50 $\pm$ 7.49 -11.83 $\pm$ 7.64 -6.32 $\pm$ 6.73
D-KF [16] -23.68 $\pm$ 6.44 -17.36 $\pm$ 4.65 -6.61 $\pm$ 4.92
@@
KF [10] -22.55 $\pm$ 6.46 -14.14 $\pm$ 5.27 8.05 $\pm$ 0.30
KNet [17] -23.19 $\pm$ 5.54 -17.34 $\pm$ 6.12 -9.81 $\pm$ 8.92
D-KF [16] -29.67 $\pm$ 6.90 -22.97 $\pm$ 5.59 -10.59 $\pm$ 7.26
@@
Quantitative Analysis. Table III reports the results across tracking frequencies (2 - 10 Hz), prediction horizons ($\mathcal{H}_{t}=5,10$), and driving scenarios (Highway, City). Across all settings, D-KNet consistently achieves the lowest NMSE and significantly outperforms all baselines. Notably, at $\mathcal{H}_{t}=5$, D-KNet attains $45.00\pm 7.35$ dB in the City scenario at 10 Hz and surpasses the second best D-KF [16] by $12.26$ dB. Even at $\mathcal{H}_{t}=10$, D-KNet maintains robustness to long-horizon uncertainty, whereas other methods suffer considerable degradation. Tighter standard deviations of D-KNet further confirm its stability across diverse obstacle dynamics. These results demonstrate the superior reliability of D-KNet, especially under real-world constraints such as varying sensor noise, limited processing frequency, and scenario complexity.
@@ V-C More Quantitative Results
We also deploy D-KNet on an NVIDIA Jetson Orin NX 16 GB to evaluate its real-time tracking efficiency on a resource-constrained platform. It is found that D-KNet consumes only about $107~\text{MB}$ of GPU memory. As Table IV presents, it achieves inference rates of over $100$ Hz when tracking a single fast-moving obstacle. Notably, even with $10$ simultaneously tracked obstacles, D-KNet sustains about $15$ Hz with low CPU usage of $42.0\,\%$. These confirm D-KNet’s suitability for lightweight deployment on robot platforms.
Table V shows per-step NMSE (dB) by mean $\pm$ std over a 10-step prediction horizon on the AevaScenes [15] dataset. D-KNet consistently outperforms all baselines at every step, with margin widening especially in short steps (e.g., 48.81-48.81 dB of DKNet and 16.49-16.49 dB of KF for step $1$ in City). The KF and D-KF methods exhibit nearly flat error growth, indicating trivial estimates under transition model mismatch. However, D-KNet and KNet show physically reasonable error accumulation, demonstrating the effectiveness of model-based learning against dynamics uncertainty. Notably, D-KNet maintains tighter standard deviations and significantly lower mean error, verifying that the Doppler LiDAR integration enhances both accuracy and robustness for long-horizon predictions.
TABLE IV: Hardware Efficiency of D-KNet
@@
(a) Precrash.
(b) Hallucination.
(c) Crash.
Figure 7: Failure mode analysis.
TABLE V: Per-horizon-step Prediction NMSE (dB) by Mean $\pm$ Std
@@
KF [10] -15.53 $\pm$ 8.80 -15.15 $\pm$ 9.63 -14.72 $\pm$ 8.45 -14.24 $\pm$ 8.26 -13.73 $\pm$ 8.09 -13.20 $\pm$ 7.90 -12.66 $\pm$ 7.72 -12.08 $\pm$ 7.53 -11.52 $\pm$ 7.33 -10.96 $\pm$ 7.15
KNet [17] -30.82 $\pm$ 3.80 -24.80 $\pm$ 4.05 -21.44 $\pm$ 4.28 -18.93 $\pm$ 4.50 -16.93 $\pm$ 4.70 -15.24 $\pm$ 4.87 -13.78 $\pm$ 5.03 -12.42 $\pm$ 5.19 -11.23 $\pm$ 5.34 -10.16 $\pm$ 5.50
@@
KF [10] -16.49 $\pm$ 7.12 -16.34 $\pm$ 6.83 -16.32 $\pm$ 6.69 -16.15 $\pm$ 6.54 -15.92 $\pm$ 6.42 -15.65 $\pm$ 6.30 -15.34 $\pm$ 6.17 -14.98 $\pm$ 6.04 -14.59 $\pm$ 5.91 -14.16 $\pm$ 5.77
KNet [17] -38.22 $\pm$ 3.05 -32.42 $\pm$ 3.10 -28.90 $\pm$ 3.16 -26.26 $\pm$ 3.42 -24.13 $\pm$ 3.57 -22.3 $\pm$ 3.72 -20.68 $\pm$ 3.88 -19.22 $\pm$ 3.96 -17.87 $\pm$ 3.95 -16.59 $\pm$ 4.06
D-KF [16] -16.29 $\pm$ 8.89 -16.26 $\pm$ 8.19 -16.23 $\pm$ 7.78 -16.19 $\pm$ 7.47 -16.17 $\pm$ 7.26 -16.13 $\pm$ 7.07 -16.13 $\pm$ 6.88 -16.13 $\pm$ 6.68 -16.12 $\pm$ 6.51 -16.11 $\pm$ 6.34
@@
V-D Limitation and Future Work
In practice, bounding box detection noises (e.g., jitters, false positives) may affect obstacle linear velocity estimation and then downstream planning. Fig. 7 demonstrates a failure mode case where a severe box under-segmentation occurs when two rapid obstacles (moving at 6.106.10\,m/s and 5.285.28\,m/s) approach, as depicted in (b). This contaminates linear velocity estimation and obstacle tracking with a hallucinated object moving at 3.453.45\,m/s and propagates to downstream planning, resulting in a collision failure eventually, shown in (c). To address this, our future work aims to develop bounding-box-free methods for Doppler LiDAR obstacle tracking. In terms of controller parameter tuning, DT-MPC explores Doppler-guided tuning with a heuristic strategy, which is in general suboptimal for minimizing $\mathcal{L}$, though proved to be efficient and effective. To address this, our future work aims to integrate gradient-informed techniques like differentiable optimization or policy search [2, 21] for optimal Doppler-guided tuning, while maintaining model efficiency.